(D^2+4D+3)y=2cos^2x


1
Expert's answer
2022-03-15T19:39:00-0400

The general solution takes the form

y=yc+yp

To solve for yc;

An auxiliary equation can be given as:

"m^2+4m+3=0"

D=16-12=4

"m_1=\frac{-4+2}{2}=-1"

"m_2=\frac{-4-2}{2}=-3"

"y_c=C_1e^{-x}+C_2e^{-3x}"

2cos2x=cos(2x)+1

"y_p=x^se^{ax}(R_m(x)cos(bx)+T_m(x)sin(bx)("

s=0 if a+bi is not a root or s=k if it is.

Sollutiong for 1:

a+bi=0, then s=0

y0=A

y'0=0

y''0=0

3A=1

A=1/3

y0=1/3


Sollution fot cos(2x)

a+bi=2i, then s=0

y1=Bsin(2x)+Acos(2x)

y'1=2Bcos(2x)-2Asin(2x)

y''1=-4Bsin(2x)-4Acos(2x)

(-B-8A)sin(2x)+(8B-A)cos(2x)=cos(2x)

A=-1/65

B=8/65

"y_1=\frac{8sin(2x)}{65}-\frac{cos(2x)}{65}"

"y=C_1e^{-x}+C_2e^{-3x}+ \frac{8sin(2x)}{65}-\frac{cos(2x)}{65}+1/3"